OpenStudy (anonymous):

how do i solve 24 4/9 take away 3 5/6

OpenStudy (lgbasallote):

take away means subtraction... \[\huge 24 \frac 49 - 3 \frac 56\] so where are you having trouble?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

yes....????

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

get the denominator be the same number so 18 would be great. so it would be 24 and 8 over 18 subtracting 3 and 15 over 18.\[24\frac{ 8 }{ 18 } -3 \frac{ 15 }{ 18 }\] \[Answer: 20 \frac{ 11 }{ 18 }\]

OpenStudy (anonymous):

if you're wondering how i got the 8 and 15 as the numerator, i multiply the top by 2 and the other fraction by 3 so the demonator will be the same.